OpenStudy (anonymous):

What is the value of x? x/4 - 5 = 6

OpenStudy (anonymous):

are you familiar with PEMDAS?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

okay (: so first, you would multiply 6 by 4, then add 5 to that, and you have your answer!

OpenStudy (anonymous):

29?

OpenStudy (anonymous):

correct!

OpenStudy (aum):

Is the problem \[ \Large \frac x4 - 5 = 6 \]?

OpenStudy (anonymous):

yeah

OpenStudy (aum):

x / 4 - 5 = 6 To solve for x, first isolate x. To do that add 5 to both sides: x / 4 = 6 + 5 x / 4 = 11 multiply both sides by 4 x = 44

OpenStudy (aum):

And you can quickly verify the answer x = 44 is correct: 44/4 - 5 = 11 - 5 = 6

OpenStudy (anonymous):

^but wouldnt you have to multiply before you add?

OpenStudy (aum):

No. PEMDAS does not apply here. The problem here is not the order of operations. It is an equation where we have to solve for x. To solve for x we have to isolate x first. Step 1. Add 5 to both sides. Step 2. Multiply both sides by 4.

OpenStudy (anonymous):

oh, well thanks for clearing it up for me! @aum

OpenStudy (aum):

You are welcome.